write the slope-y-intercept equation form of the line which has slope 2/3 and passes through the point: (-3,3)


write the point-slope equation form of the line which passes through 2 points: (-9,0), (-5,3)

write the general equation form of the line which passes through the 2 points:(1,2), (6,7)

Answer by venugopalramana(3286)

write the slope-y-intercept equation form of the line which has slope 2/3 =M SAY and passes through the point:P = (X1,Y1)SAY (-3,3)
FORMULA IS
Y= MX -MX1+Y1
Y= 2X/3 -(2/3)(-3)+3=2X/3 +5
write the point-slope equation form of the line which passes through 2 points: P= (X1,Y1) SAY=(-9,0), Q = (X2,Y2) SAY= (-5,3)
FORMULA IS
Y = (Y2-Y1)(X-X1)/(X2-X1) +Y1=(3-0)(X+9)/(-5-0) + 0
Y= -3X/5 -27/5

write the general equation form of the line which passes through the 2 points:(1,2), (6,7)
SAME AS ABOVE